OpenStudy (icalibear):

Please help, I need to write a two-column proof? Given: 7y = 8x - 14; y = 6 Prove: x = 7 @Hero

hero (hero):

Input x = 7 and y = 6 into the equation

OpenStudy (icalibear):

7(6) = 8(7)

OpenStudy (icalibear):

but I need to write the steps and stuff

OpenStudy (anonymous):

Well. You can start it off by working it out like this: 7y = 8x - 14; y = 6 So, 7(6) = 8x - 14 42 = 8x - 14 42 = 8(7) - 14 42 = 56 - 14 Then, 56 - 14 = 42. So, 42 = 42 That's your proof. :)

OpenStudy (anonymous):

Replace x for a number like finding out a simple variable by trying different numbers, but 7 it your answer because it fits into the equation.

OpenStudy (icalibear):

So like if I was going to write it out 7y = 8x - 14 ; y = 6 Given 7(6) = 8x - 14 Substitution Property of Equality 42 = 8x - 14 What would the next step be?

OpenStudy (anonymous):

Plug in x for a number until you get the correct one, which is 7. It's sorta similar to finding an answer for a simple equation with a variable in it, trying different numbers until you have the correct one.

OpenStudy (icalibear):

It's asking me for the properties though, see like the substitution property, I just don't know what property that would represent

OpenStudy (anonymous):

But here is another way of doing it, I'll start off from when you asked what the next step would be.

OpenStudy (anonymous):

42 = 8x - 14 Add 14 to both sides. 42 (+14) = 8x - 14 (+14) 56 = 8x Now plug in x with a number. x = 6 won't work because it'll equal 48. But x = 8 won't work either because it'll equal 64. x = 7 because it fits, and it finishes the equation: 56 = 8(7) 56 = 56

OpenStudy (anonymous):

You have 2 sets of proof there. :)

OpenStudy (icalibear):

So for the next step is that the addition property

OpenStudy (icalibear):

o.o

OpenStudy (anonymous):

Uh. o.o What do you mean? :u

OpenStudy (icalibear):

i know how to solve the proof but i need to figure out the properties lol its ok i can probably find it in the textbook :P thanks for helping though i wouldnt have solved it right

OpenStudy (anonymous):

Okay, lol. You're welcome. :D

OpenStudy (anonymous):

Yay. We fanned each other. xD

OpenStudy (icalibear):

xP